OpenStudy (anonymous):

2 years, 5% interest rate compounded semiannually, end goal is $17,500.00. how much does the initial investment need to be?

OpenStudy (anonymous):

N=2*2(because semi annually)=4, I-5/2=2.5%(Semi annually), FV= 17500, PMT=0, PV=? Formula, Fv= pv(1+I)^n Therefore, Pv=FV/(1+I)^N =17500/(1.025)4, =17500/1.103812891 =15854.13628 =15854 is the initial amount need to be invested.